# A Theoretical Framework for Iteratively Enhanced Room-Temperature Single-Photon Detection

### v3

@@ Abstract
High-performance photon detection is indispensable in a wide range of quantum-optical applications and is conventionally treated as a fixed device-level operation based on single-photon detectors (SPDs). However, state-of-the-art SPDs rely on superconducting materials, which impose severe technological demands and require challenging operational conditions such as cryogenic cooling, thereby hindering scalable implementation. Here, we propose the enhanced single-photon detector (ESPD) framework, a theoretical paradigm supported by numerical simulations, that reformulates single-photon detection as an iteratively enhanced process based on state preparation, controlled operations, projective measurements, and multi-copy analysis, and enables substantial performance improvement using exclusively room-temperature components. Numerical simulations indicate that, within a physically motivated parameter regime, the ESPD framework can upgrade a legacy SPD with a detection efficiency (DE) of about 59% and a dark count rate (DCR) of about $10^{-2}$ to effective performance metrics with DE exceeding 93% and DCR below $10^{-9}$, comparable to those of superconducting SPDs. As a consequence, the minimal tolerable channel transmission rate for quantum key distribution protocols can be reduced by several orders of magnitude. While its physical realization would require substantial experimental integration, the ESPD framework establishes a general system-level perspective on photon detection, highlighting the potential of iterative quantum processing for overcoming intrinsic detector limitations at room temperature.
preprint: APS/123-QED
I Introduction
Single-photon detection plays a central role in quantum optics and underlies a wide range of quantum-optical applications. Conventionally, the detection process is treated as a fixed device-level operation implemented by single-photon detectors (SPDs). The two most critical performance metrics of an SPD are the detection efficiency (DE), denoted by $\eta$, and the dark count rate (DCR), denoted by $d$. The former affects the efficiency of nearly all quantum operations, while the latter imposes the ultimate limitation in high-attenuation scenarios. In terms of DE, numerous quantum experiments rely on SPD signals to retrieve results. In linear optical quantum computation, where success events are probabilistic and require the simultaneous detection of multiple photons, the success rate decreases exponentially with low-efficiency SPDs [1, 2, 3]. In Bell tests for quantum non-locality, verifying violations of the Clauser–Horne–Shimony–Holt or Eberhard inequalities routinely requires DE exceeding specific thresholds to close loopholes [4, 5, 6, 7, 8, 9, 10], such as approximately 67% [11] or 83% [12]. In heralded single-photon sources, where entangled pairs are generated, and one photon is detected to confirm the presence of the other, the efficiency of the SPD directly dictates the success rate [13, 14, 15, 16]. With respect to DCR, a primary constraint arises in quantum communication, particularly in quantum key distribution (QKD). The security of a practical QKD protocol relies on maintaining the quantum bit error rate (QBER) below a protocol-specific threshold, significantly lower than 50% [17, 18, 19, 20]. However, communication channels induce exponential photon attenuation [21, 22, 23, 24]. If the effective signal rate, determined by the channel transmission rate and thus communication distance, falls below the DCR of the SPDs, the QBER rises precipitously, as dark count events (which yield a 50% error rate) dominate the detection statistics. Consequently, the practical secure distance of QKD is strictly limited by the DCR of the SPDs [25, 26, 27].
Driven by their critical importance, SPD has remained a continuous research focus since the inception of quantum technologies. While capable of operating at room temperature, traditional SPDs often suffer from both low DE and high DCR. For instance, a frequency upconversion SPD presented in 2004 operates at 300 K with DE about 59% and DCR of $10^{-2}$ [28], while an InGaAs/InP SPD presented in 2017 operates at 223 K with DE about 27.5% and DCR of $10^{-6}$ [29]. Such specifications are insufficient for advanced quantum tasks, which typically require the DE of SPD above 90% with DCR below $10^{-7}$.
To suppress DCR and enhance DE, recent efforts have shifted toward designing SPDs using superconducting devices, such as superconducting nanowire SPDs (SNSPDs) [30, 31, 32, 33]. These SPDs achieve performance superior to traditional counterparts, boosting DE to over 90% while suppressing DCR by several orders of magnitude. This capability enables the implementation of demanding quantum tasks, such as long-distance QKD [19, 34, 35, 36]. A summary of recent representative advancements in superconducting SPDs is presented in Table 1.
@@ I Introduction
Table 1: Summary of recent representative performance benchmarks for state-of-the-art superconducting SPDs. OT: Operation Temperature, DE: Detective efficiency, DCR: Dark count rate.
Year Researchers (et al.) OT DE DCR
@@
However, as illustrated in the table, the reliance on superconductors necessitates stringent operating conditions, specifically with extremely low temperatures. When these demanding conditions are relaxed, the SPD performance degrades significantly. This requirement introduces significant infrastructural requirements and operational complexity, fundamentally hindering the scalability and widespread implementation of high-performance quantum technology in field-deployable and environmentally constrained situations.
To address the limitations imposed by the suboptimal performance of room-temperature SPDs and cryogenic operation in superconducting alternatives, we propose the enhanced single-photon detector (ESPD) framework, a system-level theoretical paradigm supported by numerical simulations for improving single-photon detection using exclusively room-temperature components. The ESPD framework reformulates single-photon detection as an iteratively enhanced process rather than a fixed device-level operation, in which detection performance is progressively improved through the integration of state preparation, controlled operations, projective measurements, and multi-copy analysis. As a representative numerical benchmark, we show that, within a physically motivated parameter regime, the ESPD framework can enhance a legacy non-superconducting SPD reported two decades ago [28], upgrading an initial detector with DE of about 59% and DCR of $10^{-2}$ to effective performance metrics with DE exceeding 93% and DCR below $10^{-9}$. These values are comparable to those of state-of-the-art superconducting SPDs and relax the minimal tolerable channel transmission rate for quantum key distribution by several orders of magnitude. While the physical realization of such an architecture would require substantial experimental integration, the ESPD framework provides a general platform for the theoretical and numerical investigation of enhancement mechanisms in room-temperature photon detection, highlighting the way in which system-level quantum processing can overcome intrinsic device limitations.
II Method and Result
In this section, we detail the architecture of the Enhanced Single-Photon Detector (ESPD) scheme, which iteratively improves performance initialized from a base (existing) SPD. We establish the theoretical framework governing its operation, provide analytical approximations to characterize its performance, and state its applications in QKD.
II.1 Paradigm of the Enhanced Single-Photon Detection
The ESPD architecture is designed to recursively upgrade the capabilities of a given SPD level by level, ultimately achieving high DE and low DCR.
For clarity, we clarify that DE is the probability of generating a positive detection event given a non-vacuum (single-photon) input, and the DCR is the probability of a positive report given a vacuum input. Let $\text{ESPD}_{s}$ denote the ESPD at level $s$, with DE $\eta_{s}$ and DCR $d_{s}$. The iteration begins with an initial existing SPD, $\text{ESPD}_{0}$, characterized by initial value $\eta_{0}=\eta$ and $d_{0}=d$.
@@ II.1 Paradigm of the Enhanced Single-Photon Detection
The construction of the $(s+1)$-th level detector, $\text{ESPD}_{s+1}$, proceeds as follows. To facilitate understanding, readers can consider an implementation where the degree of freedom (DOF) is photon polarization and the controlled operations are C-NOT gates.
@@
    DOF Selection: Select a manipulable DOF of the incoming photon (e.g., polarization). Let $\{|0\rangle,|1\rangle\}$ denote the computational basis for this DOF. We treat the signal as a qutrit in the basis $\{|\emptyset\rangle,|0\rangle,|1\rangle\}$, where $|\emptyset\rangle$ denotes the vacuum state, and $|0\rangle,|1\rangle$ denote the orthonormal states of a non-vacuum photon [^1] Note that if the selected DOF is the photon number itself, $|0\rangle$ coincides with $|\emptyset\rangle$ and $|1\rangle$ represents non-vacuum, which degenerates to the qubit case. However, for practical implementations, it is often more convenient to use an internal DOF like polarization distinct from photon existence..
@@
    State Preparation: Apply a state preparation operation to the incoming signal to initialize the selected DOF to $|1\rangle$. Consequently, a non-vacuum signal is mapped to $|1\rangle$, while a vacuum signal remains $|\emptyset\rangle$.
@@
    Controlled Operations: Implement $n_{s+1}$ controlled operations, denoted as $C(F_{i})$ for $i=1,\dots,n_{s+1}$ (typically C-NOT gates). These gates interact the signal path with $n_{s+1}$ auxiliary sources, each initialized in the state $|0\rangle$.
@@
    Auxiliary Measurement: Perform projective measurements on the auxiliary outputs in the basis $\{|0\rangle,|1\rangle\}$. Outcome paths of $|0\rangle$ are discarded, and only the $|1\rangle$ paths are preserved for detection.
@@
    Detection: Detect the signals at the end of paths, including the signal one and the auxiliary ones, using $s$-th level detectors, $\text{ESPD}_{s}$.
@@
    Decision Logic: A positive detection report for $\text{ESPD}_{s+1}$ is generated if and only if at least $k_{s+1}$ positive detections are registered among the $n_{s+1}+1$ detectors (comprising $n_{s+1}$ auxiliary detectors and one signal path detector). Otherwise, a negative report is returned. Here, $k_{s+1}$ is a pre-selected threshold integer, and $n_{s+1}$ refers to the number of effective operations [^2] For example, if the controlled gates are deterministic, then $n_{k+1}$ is exactly the number of used operations, while if the controlled gates are probabilistic, then $n_{k+1}$ refers to the successful ones after post-selection. Therefore, the decision logic is essentially that the positive report rate (among all effective reports) on the multiple detections reaches a pre-defined threshold..
@@
Physical intuition: Intuitively, each level of the ESPD framework increases the number of effective detection opportunities for an input signal, thereby enhancing the overall detection capability. In an idealized limit where all components except the SPDs are assumed to be perfect, the $(s+1)$-th level generates $n_{s+1}$ identical copies of the incoming signal by the controlled modules, which are independently detected by $n_{s+1}$ SPDs. A non-vacuum outcome is reported if at least $k_{s+1}$ positive detection events are registered. Under this idealized assumption, the detection efficiency can be enhanced by choosing a small threshold $k_{s+1}$, for example, when $k_{s+1}=1$, the effective DE becomes $1-(1-\eta_{s})^{n_{s+1}}$. Furthermore, the DCR is strongly suppressed, since a false positive outcome occurs only when dark-count events arise in at least $k_{s+1}$ detectors, leading to an approximate scaling $d_{s}^{k_{s+1}}$, which decreases dramatically when $k_{s+1}$ is increased. This physical intuition remains valid beyond the idealized limit. When imperfections in the auxiliary components introduce errors that are sufficiently small relative to the activation probability of a genuine input signal, the ESPD framework continues to improve detection performance. A rigorous formulation of this behavior, accounting for non-ideal devices and error propagation, is developed in Section II.2.
II.2 Theoretical Formulations
@@
The rigorous theoretical formulation of the ESPD framework is developed in this subsection. We first introduce the notations and assumptions, and then present the exact analytical expressions.
II.2.1 Notations and Basic Assumptions
@@
We define the parameters governing the system dynamics. As stated, $\eta_{s}$ and $d_{s}$ represent the DE and DCR of $\text{ESPD}_{s}$, respectively. Although $\eta_{s}$ may depend on the photon number distribution, we approximate it as a unified value for the given system for simplicity. The parameters $n_{s+1}$ and $k_{s+1}$ denote the number of effective controlled operations and the decision threshold for $\text{ESPD}_{s+1}$, respectively.
@@
For convenience, assume that all controlled operations $C(F_{i})$ are identical, denoted simply as $C(F)$, and they are implemented in a feedforward controlled configuration, which ensures that $n_{s+1}$ represents both the number of effective modules and the total physical modules required for level $s+1$. However, we emphasize that the generalized ESPD paradigm remains valid for post-selection controlled gates (for a brief discussion, see Section III.1.2). Let $p$ be the transmission probability of the signal path through a single module (an (optional) state preparation followed by a $C(F)$). Specifically, $p$ is the probability that the output signal remains non-vacuum given a non-vacuum input of a single non-cascaded module, accounting for loss but irrespective of decoherence (since the state can be re-prepared). Let $P$ and $Q$ denote the probabilities of obtaining a non-vacuum output on the controlled path before detection, given that the input to the corresponding module was non-vacuum or vacuum, respectively. Hence, $Q$ represents the intrinsic error rate of the auxiliary path in the absence of $C(F)$ (i.e., errors from auxiliary state preparation and measurement (SPAM), but not including detection). We approximately assume $p$, $P$, and $Q$ are invariant across all $C(F)$ module operations.
@@
II.2.2 Parameter Approximation Assumptions
The assumptions below are employed solely to derive analytical approximations in Section II.3, as well as the approximation part in Eq. (1) and (2). They are not required for the general theoretical formulation of the ESPD framework nor for the numerical simulations presented in this work. Although these assumptions entail relatively stringent requirements, they are physically motivated by realistic components and will be discussed in detail in Section III.1.
In practical scenarios, SPD DE typically ranges from 10% to 90%, while the DCR is generally below $10^{-3}$. Consequently, the hierarchy $0\leq d\ll\eta\leq 1$ holds. We further assume $d_{s}\ll P\eta_{s}$, which is trivial: If it were violated, the signal contribution from the controlled operations would become indistinguishable from device noise, rendering the operations practically insignificant. Additionally, we assume that the transmission rate of a single controlled module $p\approx 1$, SPAM error of the auxiliary path $Q\ll 1$, given the maturity of commercial optical components (see Section III.1). Finally, considering practical implementation costs, we assume $n_{s}$ and $k_{s}$ are moderate integers satisfying $1\leq k_{s}\leq n_{s}$.
II.2.3 Formulations
Let $P_{s}$ and $Q_{s}$ denote the probabilities of obtaining a positive report on an auxiliary path given a non-vacuum or vacuum input to a $C(F)$ module, respectively, using $\text{ESPD}_{s}$ as the detector. They can be expressed as:
$P_{s+1}$ $=P[\eta_{s}+(1-\eta_{s})d_{s}]+(1-P)d_{s}$ (1)
$=P\eta_{s}(1-d_{s})+d_{s}\approx P\eta_{s}$
$Q_{s+1}$ $=Q[\eta_{s}+(1-\eta_{s})d_{s}]+(1-Q)d_{s}$ (2)
$=Q\eta_{s}(1-d_{s})+d_{s}\approx Q\eta_{s}+d_{s}\leq Q+d_{s}$
Similarly, let $P^{\prime}_{s}$ and $Q^{\prime}_{s}$ denote the probabilities of obtaining a positive report on the signal path after all $C(F)$ operations, given a non-vacuum or vacuum output from the final controlled gate, respectively, then:

### v4

# Beyond Single-Device Constraints: A System-Level Theoretical Framework for High-Performance Single-Photon Detection at Room Temperature

@@ Abstract
Photon detection, a fundamental quantum technology, is traditionally treated as a fixed device-level operation governed by intrinsic properties of single-photon detectors (SPDs). High-performance detection has therefore largely relied on superconducting technologies whose requirement for cryogenic operation imposes substantial infrastructure constraints, limiting scalable deployments. Here, the enhanced single-photon detection (ESPD) framework is presented as a system-level theoretical paradigm that shifts photon detection from device-centric optimization to an integrated quantum-information-processing task, by reformulating it as an iteratively enhanced process integrating state-preparation, controlled operations, projective measurements, and multi-copy analysis. ESPD enables systematic performance enhancement through architectural design rather than material modification, thereby circumventing superconducting components, allowing high-performance detection using exclusively room-temperature hardware. Numerical simulations grounded in physically motivated parameters indicate that the ESPD framework can upgrade a conventional room-temperature SPD to effective DE exceeding 93% and DCR below $10^{-9}$, which are comparable to state-of-the-art superconducting SPDs and can significantly relax the minimal tolerable channel transmission rate in quantum communications. While physical realization would require further component integration, this work establishes a rigorous theoretical framework for enhancing detection performance through architectural quantum-information principles, providing a general blueprint for transcending device-level constraints and guiding the development of next-generation room-temperature quantum technologies.
keywords:
Single-photon detection, System-level quantum architecture, Iterative quantum information processing, Room-temperature quantum technologies, Photonic integration
Hao Shu*
[…]
{affiliations}
Dr. Hao Shu
Address: Sun Yat-Sen University, Guangzhou, China
Email Address: Hao_B_Shu@163.com (Primary), shuh6@mail.sysu.edu.cn (Institutional)
1 Introduction
Single-photon detection plays a central role in quantum optics and serves as the bedrock for a wide range of quantum-information applications. Conventionally, the detection process is treated as a fixed device-level operation primarily determined by the material properties of single-photon detectors (SPDs). The primary performance metrics characterizing an SPD are the detection efficiency (DE), denoted by $\eta$, and the dark count rate (DCR), denoted by $d$. The former directly dictates the success probability of quantum operations conditioned on detection events, while the latter imposes a fundamental signal-to-noise constraint in high-attenuation scenarios.
In terms of DE, high-performance detection is indispensable for retrieving meaningful outcomes in quantum experiments. In linear optical quantum computation, where success events are probabilistic and require the simultaneous detection of multiple photons, the overall success probability decreases exponentially with imperfect DE [1, 2, 3]. In fundamental tests of quantum nonlocality, verifying violations of the Clauser-Horne-Shimony-Holt or Eberhard inequalities requires DE exceeding strict thresholds to close detection loopholes [4, 5, 6, 7, 8, 9, 10], such as approximately 67% [11] or 83% [12]. In heralded single-photon sources, where one photon of an entangled pair signals the presence of the other, the achievable heralding rate is also strictly limited by DE [13, 14, 15, 16].
With respect to DCR, a particularly stringent constraint arises in quantum communication, notably in quantum key distribution (QKD). The security of practical QKD protocols relies on maintaining the quantum bit error rate (QBER) below protocol-dependent thresholds that are significantly lower than 50% [17, 18, 19, 20]. However, optical communication channels inevitably induce exponential photon attenuation with distance [21, 22, 23, 24]. When the effective signal rate, determined by channel transmission for a given DE, approaches the DCR floor of the SPDs, detection events become dominated by dark counts, which contribute a 50% error rate, causing the QBER to rise precipitously. Consequently, the achievable secure distance of QKD systems is ultimately bounded by the DCR of the employed SPDs [25, 26, 27].
Motivated by these critical demands, the development of high-performance SPDs has remained a primary research focus. While conventional semiconductor-based SPDs offer the advantage of non-cryonic operation, they typically suffer from an intrinsic trade-off between limited DE and relatively high DCR. For instance, a frequency upconversion SPD demonstrated in 2004 operated at 300 K with a DE of approximately 59% and a DCR of $10^{-2}$ [28], while an InGaAs/InP SPD reported in 2017 operated at 223 K achieved a DE of about 27.5% with a DCR of $10^{-6}$ [29]. Such specifications often fall short of the requirements for advanced quantum protocols, which frequently demand DE exceeding 90% and DCR below $10^{-7}$.
To meet these stringent requirements, substantial effort has shifted toward designing SPDs with superconducting technologies, most notably superconducting nanowire SPDs (SNSPDs) [30, 31, 32, 33]. These devices have successfully pushed DE beyond 90% while suppressing DCR by several orders of magnitude compared to semiconductor counterparts, thereby enabling demanding applications such as long-distance QKD [19, 34, 35, 36]. Recent representative benchmarks of state-of-the-art superconducting SPDs are summarized in Table 1.
@@ 1 Introduction
Table 1: Summary of recent representative performance benchmarks for state-of-the-art superconducting SPDs.
Year Researchers (et al.) Operation Temperature Detection Efficiency Dark Count Rate
@@
However, the superior performance of superconducting SPDs comes at the cost of cryogenic operation. This imposes stringent environmental requirements and necessitates bulky, power-intensive cooling infrastructure. The significant degradation of performance at elevated temperatures highlights a strong trade-off between detection performance and operating conditions. These constraints present major barriers to scalable deployment, particularly in field applications, satellite payloads, and scenarios with strict size, weight, and power limitations where cryogenic infrastructure is impractical.
To address the limitations of room-temperature SPDs without resorting to cryogenic conditions, the enhanced single-photon detection (ESPD) framework is introduced, shifting the focus from device-centric optimization to a system-level architectural paradigm. Within this framework, single-photon detection is reformulated as a composite quantum-information-processing task that integrates state preparation, controlled quantum operations, projective measurements, and multi-copy decision analysis, rather than relying solely on the material properties of a single SPD. Through structured system-level integration, the ESPD framework enables effective enhancement of detection performance using exclusively room-temperature components.
Numerical simulations, grounded in physically motivated parameters drawn from commercially available devices and experimentally reported SPDs, indicate that, under an exclusively room-temperature component setting, the ESPD framework enables effective DE exceeding 93% together with effective DCR below $10^{-9}$, starting from a legacy SPD (for example, the one reported two decades ago with about 59% DE and $10^{-2}$ DCR [28]). These effective performance metrics are comparable to those of state-of-the-art superconducting SPDs and lead to a substantial relaxation of the minimal tolerable channel transmission rate in QKD.
While physical implementation would require further system-level integration of quantum technologies, the ESPD framework provides a structured theoretical basis for achieving high-performance photon detection at room temperature. By explicitly analyzing technological feasibility and the trade-off between resource overhead and operating temperature, the ESPD framework offers a feasible technological roadmap and practical guidance for future experimental efforts. It serves as a theoretical blueprint for overcoming intrinsic device limitations, with the potential to unlock a wide range of quantum applications in environmentally constrained scenarios.
2 Method and Result
This section details the architectural structure of the ESPD framework and formulates the corresponding detection process. The ESPD paradigm is introduced in Section 2.1, with theoretical formulations established in Section 2.2. Analytical approximations are developed in Section 2.3, and numerical simulations are presented in Section 2.4. Finally, implications for QKD are discussed in Section 2.5.
2.1 Paradigm of the Enhanced Single-Photon Detection
The ESPD architecture is formulated as a level-by-level framework in which the effective detection performance of a given SPD is enhanced through successive processing stages. Here, the enhancement is quantified in terms of DE and DCR, defined respectively as the probability of generating a positive detection event given a non-vacuum (single-photon) input and the probability of a positive report given a vacuum input.
Let $\text{ESPD}_{s}$ denote the ESPD at level $s$, with effective DE $\eta_{s}$ and effective DCR $d_{s}$. The iteration begins with an existing SPD, denoted as $\text{ESPD}_{0}$, characterized by the initial values $\eta_{0}=\eta$ and $d_{0}=d$.
@@ 2.1 Paradigm of the Enhanced Single-Photon Detection
The construction of the $(s+1)$-th level detector, $\text{ESPD}_{s+1}$, is defined as follows. As a representative example, an implementation may be considered in which the degree of freedom (DOF) is photon polarization and the controlled operations correspond to controlled-NOT (C-NOT) gates.
@@
    DOF Selection: A manipulable DOF of the incoming photon is chosen (for example, polarization). The computational basis for this DOF is denoted by $\{|0\rangle,|1\rangle\}$. The signal is modeled as a qutrit in the basis $\{|\emptyset\rangle,|0\rangle,|1\rangle\}$, where $|\emptyset\rangle$ denotes the vacuum state, and $|0\rangle$ and $|1\rangle$ denote orthonormal states of a non-vacuum photon [^1] If the selected DOF is the photon number itself, $|0\rangle$ coincides with $|\emptyset\rangle$ and $|1\rangle$ represents a non-vacuum state, reducing the description to qubit. In practical architectures, an internal DOF such as polarization, distinct from photon existence, is often more convenient..
@@
    State Preparation: A state-preparation operation is applied to initialize the selected DOF to $|1\rangle$. As a result, a non-vacuum signal is mapped to $|1\rangle$, while a vacuum input remains in state $|\emptyset\rangle$.
@@
    Controlled Operations: A set of $n_{s+1}$ controlled operations, denoted as $C(F_{i})$ for $i=1,\dots,n_{s+1}$ (typically C-NOT gates), are applied to couple the signal path with $n_{s+1}$ auxiliary photons, each initialized in state $|0\rangle$.
@@
    Auxiliary Measurement: Projective measurements are performed on the auxiliary outputs in basis $\{|0\rangle,|1\rangle\}$. Outcome paths corresponding to $|0\rangle$ are discarded, and only the $|1\rangle$ outcomes are retained for subsequent detection.
@@
    Detection: Detection is carried out on all remaining paths, including both the signal path and the auxiliary paths, using $s$-th level detectors ($\text{ESPD}_{s}$).
@@
    Decision Logic: A positive detection report for $\text{ESPD}_{s+1}$ is generated if and only if at least $k_{s+1}$ positive detection events are registered among the $n_{s+1}+1$ detectors, consisting of $n_{s+1}$ auxiliary detectors and one detector on the signal path. Otherwise, a negative report is returned. Here, $k_{s+1}$ is a preselected threshold integer, and $n_{s+1}$ denotes the number of effective controlled operations [^2] The main discussions of the article assume deterministic controlled gates, for which $n_{s+1}$ equals the number of implemented operations. For probabilistic gates, $n_{s+1}$ refers to the number of successful operations after post-selection. In this case, the decision logic is naturally reformulated in terms of the fraction of positive detection outcomes among the effective detections exceeding a predefined threshold..
@@
Physical intuition: From a detection perspective, each level of the ESPD framework increases the number of effective detection opportunities available to an input signal. In an idealized limit where all components except the SPDs are assumed to be perfect, the $(s+1)$-th level produces $n_{s+1}$ effective signal instances through the controlled modules, which are independently detected by $n_{s+1}$ SPDs. A non-vacuum outcome is reported if at least $k_{s+1}$ positive detection events are registered. Under this idealized assumption, the effective DE increases with decreasing threshold $k_{s+1}$; for instance, when $k_{s+1}=1$, the effective DE is given by $1-(1-\eta_{s})^{n_{s+1}}$. Furthermore, the effective DCR is suppressed, since a false positive outcome requires dark-count events to occur in at least $k_{s+1}$ detectors, leading to an approximate scaling $d_{s}^{k_{s+1}}$. This physical intuition extends beyond the idealized limit when imperfections in the auxiliary components introduce errors that remain small compared to the activation probability of a genuine input signal. A rigorous formulation of this behavior, accounting for non-ideal devices and error propagation, is developed in Section 2.2.
2.2 Theoretical Formulations
@@
The rigorous theoretical formulation of the ESPD framework is developed in this subsection. The relevant notations and assumptions are introduced first, followed by the exact analytical expressions and the approximation analysis.
2.2.1 Notations and Basic Assumptions
The parameters governing the system behavior are defined as follows.
@@ 2.2.1 Notations and Basic Assumptions
As stated, $\eta_{s}$ and $d_{s}$ represent the DE and DCR of $\text{ESPD}_{s}$, respectively. Although $\eta_{s}$ may depend on the photon number distribution, it is approximated as a unified value for the given system for simplicity. The parameters $n_{s+1}$ and $k_{s+1}$ denote the number of effective controlled operations and the decision threshold for $\text{ESPD}_{s+1}$, respectively.
@@
For convenience, all controlled operations $C(F_{i})$ are assumed to be identical and are denoted simply as $C(F)$. They are implemented in a feedforward controlled configuration, such that $n_{s+1}$ represents both the number of effective modules and the total number of physical modules required at level $s+1$. However, the generalized ESPD paradigm remains valid for post-selection controlled gates (see Section 3.1.2 for a brief discussion). Let $p$ be the transmission probability of the signal path through a single module (an (optional) state preparation followed by a $C(F)$). Specifically, $p$ is the probability that the output signal remains non-vacuum given a non-vacuum input of a single non-cascaded module, accounting for loss but irrespective of decoherence (since the state can be re-prepared). Let $P$ and $Q$ denote the probabilities of obtaining a non-vacuum output on the controlled path before detection, given that the input to the corresponding module was non-vacuum or vacuum, respectively. Hence, $Q$ represents the intrinsic error rate of the auxiliary path in the absence of $C(F)$ (i.e., errors from auxiliary state preparation and measurement (SPAM), but not including detection). The parameters $p$, $P$, and $Q$ are approximately assumed to be invariant across all $C(F)$ module operations.
@@
2.2.2 Parameter Approximation Assumptions
The assumptions below are employed solely to derive analytical approximations in Section 2.3, as well as the approximation part in Eq. (1) to (3). They are not required for the general theoretical formulation of the ESPD framework nor for the numerical simulations presented in this work. Although these assumptions entail relatively stringent requirements, they are physically motivated by realistic components and will be discussed in detail in Section 3.1.
In practical scenarios, SPD DE typically ranges from 10% to 90%, while the DCR is generally below $10^{-3}$. Consequently, the hierarchy $0\leq d\ll\eta\leq 1$ holds. It is further assumed that $d_{s}\ll P\eta_{s}$: If this condition were violated, the signal contribution from the controlled operations would become indistinguishable from device noise, rendering the operations practically insignificant. Additionally, the transmission rate of a single controlled module is assumed to satisfy $p\approx 1$, and the SPAM error of the auxiliary path is assumed to satisfy $Q\ll 1$, consistent with the maturity of commercial optical components (see Section 3.1). Finally, considering practical implementation costs, $n_{s}$ and $k_{s}$ are taken to be moderate integers satisfying $1\leq k_{s}\leq n_{s}$.
2.2.3 Formulations
Let $P_{s}$ and $Q_{s}$ denote the probabilities of obtaining a positive report on an auxiliary path given a non-vacuum and vacuum input to a $C(F)$ module, respectively, using $\text{ESPD}_{s}$ as the detector. They can be expressed as:
$P_{s+1}=P[\eta_{s}+(1-\eta_{s})d_{s}]+(1-P)d_{s}=P\eta_{s}(1-d_{s})+d_{s}\approx P\eta_{s}$ (1)
$Q_{s+1}$ $=Q[\eta_{s}+(1-\eta_{s})d_{s}]+(1-Q)d_{s}=Q\eta_{s}(1-d_{s})+d_{s}\approx Q\eta_{s}+d_{s}\leq Q+d_{s}$ (2)
Similarly, let $P^{\prime}_{s}$ and $Q^{\prime}_{s}$ denote the probabilities of obtaining a positive report on the signal path after all $C(F)$ operations, given a non-vacuum and vacuum output from the final controlled gate, respectively, then: